1/1

William Wordsworth,  William Wordsworth,
The Daffodils  Die Narzissen

tradukita de Hans Hennecke

I wander’d lonely as a cloud That floats on high o’er vales and hills, When all at once I saw a crowd, A host of golden daffodils, Beside the lake, beneath the trees Fluttering and dancing in the breeze.

Argiope-informo:

La aspekto de tiu ĉi versio estas barita.

Supozeble la aŭtoro decidis tion.

Continuous as the stars that shine And twinkle on the milky way, They stretch’d in never-ending line Along the margin of a bay: Ten thousand saw I at a glance Tossing their heads in sprightly dance.

The waves beside them danced, but they Out-did the sparkling waves in glee: - A poet could not but be gay In such a jocund company: I gazed - and gazed - but little thought What wealth the show to me had brought.

For oft, when on my couch I lie In vacant or in pensive mood, They flash upon that inward eye Which is the bliss of solitude; And then my heart with pleasure fills And dances with the daffodils.

Verkinto de tiu ĉi Angla poemo estas William Wordsworth (∗1770 – †1850).

Arg-39-73 (2003-10-17 22:02:21)

Origina anglalingva teksto de WILLIAM WORDS-WORTH (1770 - 1850)

Traduko de la Angla poemo “The Daffodils” de William Wordsworth (∗1770 – †1850) en la Germanan de Hans Hennecke.

Arg-39-75 (2004-02-01 18:36:59)

Germanigo de la origina anglalingva teksto de WILLIAM WORDSWORTH, fare de Heneko, Johano (HANS HENNECKE). (Mi trovis tiun ci poemon en la libro �Snapshots of Britain“ / �Britannien in kleinen Geschichten“, rakontita de JOY BROWNING, eldonejo Deutscher Taschenbuchverlag, ISBN 3-423-09319-6.)

http://www.poezio.net/version?poem-id=39&version-id=73,75